# A Formal Comparison Between Chain-of-Thought and Latent Thought

### v1

@@ B.3 Looped Transformer
$f\coloneq(\operatorname{id}+\mathbf{FF}_{L})\circ(\operatorname{id}+\mathbf{SA}_{L})\circ\cdots\circ(\operatorname{id}+\mathbf{FF}_{1})\circ(\operatorname{id}+\mathbf{SA}_{1}),$
@@
$f_{\mathrm{loop}}^{k+1}(x)\coloneq f\left(f_{\mathrm{loop}}^{k}(x)\right).$
@@
Appendix C Deferred Proofs for Section 3
C.1 Precision Modeling
Following the definition of finite precision in Li et al. (2024), we focus here, for simplicity, on signed floating-point numbers without exponents.
Definition C.1 (Floating-point Representation, cf. (Li et al., 2024)).
@@ C.1 Precision Modeling
Definition C.2 (Correct Rounding, cf. (Li et al., 2024)).
@@
Definition C.3 (Summation with Iterative Rounding (Li et al., 2024)).
@@
Based on this definition, all computations in the Transformer block of Definition B.1 can be represented in finite precision. The inner product and the matrix product are defined as
@@
C.2 Definition of Assumption 3.4
Our definition of polynomially-efficient approximation is based on that of Feng et al. (2023), but differs in domain. Whereas their focus is on real-valued functions, ours concerns symbolic functions.
Definition C.4 (Polynomially-efficient approximation).
We say that a function $f_{n}:\Sigma^{\ell(n)}\to\Sigma$ admits a polynomially-efficient approximation if there exists a log-precision feedforward network $\mathrm{FF}:\mathbb{F}_{s(n)}^{\ell(n)\cdot|\Sigma|}\to\mathbb{F}_{s(n)}^{|\Sigma|},\ s(n)=O(\log n),$ such that the following holds: for every input ${\bm{x}}=(x_{1},\ldots,x_{\ell(n)})\in\Sigma^{\ell(n)}$, every error tolerance $\delta>0$,
@@ C.3 Technical lemmas
C.3.1 Orthogonal Vectors
We follow the notation of Li et al. (2024). For any positive integer $s\in\mathbb{N}^{+}$ and $x\in\{0,1,\dots,2^{s}-1\}$, we denote by $\mathsf{bin}_{s}(x)\in\{0,1\}^{s}$ the standard binary representation of $x$ using $s$ bits, defined such that $x=\sum_{i=1}^{s}2^{i}\cdot(\mathsf{bin}_{s}(x))_{i}.$ We further define the signed binary encoding of $x$, denoted by $\mathsf{sbin}_{s}(x)\in\{-1,1\}^{s}$, as $\mathsf{sbin}_{s}(x)=2\cdot\mathsf{bin}_{s}(x)-(1,\ldots,1).$ Let ${\bm{x}},{\bm{y}}\in\mathbb{R}^{s}$ be two vectors of the same length. We define their interleaving, denoted by ${{\bm{x}}}^{\frown}{{\bm{y}}}\in\mathbb{R}^{2s}$, as follows: $({{\bm{x}}}^{\frown}{{\bm{y}}})_{2i-1}=x_{i},({{\bm{x}}}^{\frown}{{\bm{y}}})_{2i}={\bm{y}}_{i}$ for all $i\in[s].$ The orthogonal vectors under finite-precision arithmetic can be:
Lemma C.5 (Li et al., 2024).
@@ C.3.1 Orthogonal Vectors
C.3.2 Position Selector
Lemma C.6.
For any $m\in\mathbb{N}^{+}$ and ${\bm{x}}\in\mathbb{F}^{s}$, there exists a feedforward layer $\mathrm{FF}:\mathbb{F}^{2s}\to\mathbb{F}^{2s}$, for any $i\in[m]$, such that
@@ C.3.2 Position Selector
Let the input be ${\bm{z}}=({\bm{x}},{\bm{e}}_{i})\in\mathbb{F}^{2s}$. Set the weight ${\bm{W}}_{1}\in\mathbb{F}^{s\times 2s}$ and bias ${\bm{b}}\in\mathbb{F}^{s}$ by
${\bm{W}}_{1}=\bigl[\,{\bm{I}}_{s}\;\;\;-B_{s}{\bm{I}}_{s}\,\bigr],\qquad{\bm{b}}={\bm{0}},$ (6)
@@
Next, set the second linear layer ${\bm{W}}_{2}\in\mathbb{F}^{2s\times s}$ by
${\bm{W}}_{2}=\begin{bmatrix}-{\bm{I}}_{s}\\ \mathbf{0}_{s\times s}\end{bmatrix}$. (9)
@@
C.3.3 Feedforward layers
Lemma C.7.
@@ C.3.3 Feedforward layers
For each $i\in[N]$, by Definition C.4, there exist width $w_{i}\in\mathbb{N}$ and parameters
${\bm{W}}^{(i)}_{1}\in\mathbb{F}_{s}^{w_{i}\times\ell_{i}|\Sigma|},\quad{\bm{W}}^{(i)}_{2}\in\mathbb{F}_{s}^{|\Sigma|\times w_{i}},\quad{\bm{b}}^{(i)}\in\mathbb{F}_{s}^{w_{i}}$
@@
${\bm{W}}_{1}\coloneqq\bigoplus_{i=1}^{N}{\bm{W}}^{(i)}_{1},\qquad{\bm{W}}_{2}\coloneqq\bigoplus_{i=1}^{N}{\bm{W}}^{(i)}_{2},\qquad{\bm{b}}\coloneqq\bigoplus_{i=1}^{N}{\bm{b}}^{(i)}.$
@@
$\mathrm{FF}({\bm{x}})\coloneqq{\bm{W}}_{2}\mathrm{ReLU}({\bm{W}}_{1}{\bm{x}}+{\bm{b}})$
@@
Lemma C.8.
Let $f:\Sigma^{\ell}\to\Sigma$ be a function that admits polynomially-efficient approximation. Then, there exist two feedforward layers
$\mathrm{FF}_{1}:\mathbb{F}_{s}^{(1+\ell)|\Sigma|+1}\to\mathbb{F}_{s}^{(1+\ell)|\Sigma|+1},\quad\mathrm{FF}_{2}:\mathbb{F}_{s}^{(1+\ell)|\Sigma|+1}\to\mathbb{F}_{s}^{(1+\ell)|\Sigma|+1},$
@@
Let $n\coloneqq|\Sigma|$ and $L\coloneqq\ell n$. Write ${\bm{u}}=(\bm{e}(x_{1}),\ldots,\bm{e}(x_{\ell}))\in\{0,1\}^{L}$. By Definition C.4, there exist $w_{f}\in\mathbb{N}$, matrices ${\bm{W}}^{(f)}_{1}\in\mathbb{F}_{s}^{w_{f}\times L}$, ${\bm{W}}^{(f)}_{2}\in\mathbb{F}_{s}^{n\times w_{f}}$, and bias ${\bm{b}}^{(f)}\in\mathbb{F}_{s}^{w_{f}}$ such that
@@
${\bm{W}}^{(1)}_{1}=\begin{bmatrix}{\bm{0}}&{\bm{W}}^{(f)}_{1}&\mathbf{0}_{\,w_{f}\times 1}\\ {\bm{0}}&{\bm{I}}_{L}&\mathbf{0}_{\,L\times 1}\end{bmatrix},\quad{\bm{b}}^{(1)}=\begin{bmatrix}{\bm{b}}^{(f)}\\ {\bm{0}}_{L}\end{bmatrix},\quad{\bm{W}}^{(1)}_{2}=\begin{bmatrix}{\bm{W}}^{(f)}_{2}&\,{\bm{0}}\\ \mathbf{0}_{\,(L-n)\times w_{f}}&\,{\bm{0}}\\ \mathbf{0}_{\,1\times w_{f}}&{\bm{0}}\end{bmatrix}$, (14)
and define $\mathrm{FF}_{1}({\bm{u}},t)\coloneqq{\bm{W}}^{(1)}_{2}\,\mathrm{ReLU}\big({\bm{W}}^{(1)}_{1}({\bm{u}},t)+{\bm{b}}^{(1)}\big)$. Then, it holds that
@@
Thus we have $(\operatorname{id}+\mathrm{FF}_{1})({\bm{u}},t)=\bigl(\mathrm{FF}_{f}({\bm{u}}),\,{\bm{0}},\,t\bigr).$
@@
${\bm{W}}^{(2)}_{1}=\begin{bmatrix}2{\bm{I}}_{n}&M{\bm{1}}_{n}\\ 2{\bm{I}}_{n}&M{\bm{1}}_{n}\\ {\bm{I}}_{n}&0\\ -{\bm{I}}_{n}&0\end{bmatrix},\ {\bm{b}}^{(2)}=\begin{bmatrix}-M{\bm{1}}_{n}\\ (1-M){\bm{1}}_{n}\\ {\bm{0}}_{n}\\ {\bm{0}}_{n}\end{bmatrix},\ {\bm{W}}^{(2)}_{2}=\begin{bmatrix}\ {\bm{I}}_{n}&-{\bm{I}}_{n}&-{\bm{I}}_{n}&\ {\bm{I}}_{n}\ \\ \ \mathbf{0}_{\,(L-n)\times n}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \ \mathbf{0}_{\,1\times n}&\mathbf{0}&\mathbf{0}&\mathbf{0}\end{bmatrix}$, (18)
@@
$\mathrm{FF}_{2}({\bm{z}},{\bm{0}},t)$ $={\bm{W}}^{(2)}_{2}\,\begin{bmatrix}\mathrm{ReLU}\bigl(2{\bm{z}}+Mt{\bm{1}}_{n}-M{\bm{1}}_{n}\bigr)\\ \mathrm{ReLU}\bigl(2{\bm{z}}+Mt{\bm{1}}_{n}+(1-M){\bm{1}}_{n}\bigr)\\ \mathrm{ReLU}({\bm{z}})\\ \mathrm{ReLU}(-{\bm{z}})\end{bmatrix}$ (19)
@@
Therefore, the composition satisfies $(\operatorname{id}+\mathrm{FF}_{2})\circ(\operatorname{id}+\mathrm{FF}_{1})\bigl({\bm{u}},t\bigr)=(t\cdot\bm{e}(f({\bm{x}})),\ {\bm{0}}_{L},\ t).$
C.4 Proof for Theorem 3.5
@@
$\bm{\mathrm{sbinpred}}_{s(n)}(v)=\left(\mathsf{sbin}_{s(n)}(\mathrm{pred}(v_{i})_{0}),\,\ldots,\,\mathsf{sbin}_{s(n)}(\mathrm{pred}(v_{i})_{C_{\mathrm{max}}(n)})\right).$ (25)
@@
${\bm{q}}_{i,h}$ $={\mathsf{sbin}_{s(n)}(\mathrm{pred}(v_{i})_{h})}^{\frown}{1_{s(n)}},$ (27)
@@
where vectors of different lengths are zero-padded to match the dimension. By Lemma C.5, each attention head of the last position $i=n+k$ retrieves the predecessor’s value of $v_{n+k}$
${\bm{a}}_{n+k,h}=\bm{e}\bigl(v_{\mathrm{pred}(v_{n+k})_{h}}(x)\bigr)$ (30)
@@
${\bm{O}}({\bm{a}}_{i,1},\ldots,{\bm{a}}_{i,H})=({\bm{0}},\,\bm{e}(v_{\mathrm{pred}(v_{n+k})_{0}}(x)),\,\ldots,\,\bm{e}(v_{\mathrm{pred}(v_{n+k})_{C_{\mathrm{max}(n)}}}(x)),\,{\bm{0}}),$ (31)
@@
$\underbrace{\bm{e}(v_{\mathrm{pred}(v_{n+k})_{0}}(x)),\,\ldots,\,\bm{e}(v_{\mathrm{pred}(v_{n+k})_{C_{\mathrm{max}(n)}}}(x))}_{\text{updated}}\,,{\bm{0}}_{C_{\mathrm{sum}}(n)|\Sigma|}\Big).$ (33)
@@
By Lemma C.7, a single feed-forward layer can approximate multiple functions by partitioning the input into blocks. The first feed-forward layer then places the arguments, gathered by attention, into the correct positions. By Lemma C.6, where the vector ${\bm{e}}_{i}$ therein corresponds to ${\bm{1}}_{|\Sigma|}$ here, the hidden state at the last position, denoted by ${\bm{h}}^{(1)}_{n+k}$, becomes
$\Big(({\bm{h}}^{(0.5)}_{n+k})_{1:r},\big\|_{j=1}^{|\mathcal{F}|}\,\big(\bm{e}(v_{\mathrm{pred}(v_{n+k})_{1}}(x))\cdot 1,\ldots,\bm{e}(v_{\mathrm{pred}(v_{n+k})_{C_{j}(n)}}(x))\cdot 1\Big),$ (34)
@@
By Assumption 3.4 and Lemmas C.8 and C.7, there exist feed-forward layers $\mathrm{FF}_{2},\mathrm{FF}_{3}:\mathbb{F}_{s(n)}^{m}\to\mathbb{F}_{s(n)}^{m}$ such that, for every input tuple ${\bm{x}}_{j}\coloneq(x^{(j)}_{1},\ldots,x^{(j)}_{C_{j}(n)})\in\Sigma^{C_{j}(n)}$ for $j\in[|\mathcal{F}|]$ and every ${\bm{t}}\in\{0,1\}^{|\mathcal{F}|}$, the composition $\mathcal{FF}\coloneq(\operatorname{id}+\mathrm{FF}_{3})\circ(\operatorname{id}+\mathrm{FF}_{2})$ satisfies
@@
$=\mathcal{FF}\left(*,\,\bm{e}(v_{n+k+1}),\,*,\,\bigg\|_{j=1}^{|\mathcal{F}|}\,\big(\bm{e}(v_{\mathrm{pred}(v_{n+k})_{1}}(x)),\ldots,\bm{e}(v_{\mathrm{pred}(v_{n+k})_{C_{j}(n)}}(x))\big)\right)$ (37)
$=\Biggl(\sum_{j=1}^{|\mathcal{F}|}\bm{e}(v_{n+k+1})_{j}\cdot\bm{e}\Big(f_{j}\big(\bm{e}(v_{\mathrm{pred}(v_{n+k})_{1}}(x)),\ldots,\bm{e}(v_{\mathrm{pred}(v_{n+k})_{C_{j}(n)}}(x))\big)\Big),\;*\Biggr)$ (38)
$=\Biggl(\bm{e}\Big(f_{l}\big(\bm{e}(v_{\mathrm{pred}(v_{n+k})_{1}}(x)),\ldots,\bm{e}(v_{\mathrm{pred}(v_{n+k})_{C_{l}(n)}}(x))\big)\Big),\;*\Biggr),\,\text{where}\quad\bm{e}(v_{n+k+1})={\bm{e}}_{l}$ (39)
@@
C.5 Proof for Theorem 3.7
Figure 5: Illustration of the role of the attention and feedforward layers in Looped TFs for evaluating DAGs: the attention layer uniformly attends to and aggregates all inputs at each position, while the feedforward layer simultaneously simulates the functions of all DAG nodes in parallel.
@@
We construct a model in which (1) the attention layer aggregates the inputs, and (2) the looped feed-forward layer performs the computation of all nodes in parallel, as illustrated in Figure 5.
We first show that a feedforward layer followed by an attention layer can copy all input tokens to the each position, as shown below. Assume, given an input sequence $x=(x_{1},x_{2},\ldots,x_{n})\in\Sigma^{n}$ and one-hot encoding $\bm{e}:\Sigma\to\{0,1\}^{|\Sigma|}$. Assume each input at position $i\in[n]$ is embedded as

### v2

@@ A.4 Continuous Thought
$\mathrm{TF}_{\mathrm{dec}}\coloneq(\operatorname{id}+\mathbf{FF}_{L})\circ(\operatorname{id}+\mathbf{SA}_{L})\circ\cdots\circ(\operatorname{id}+\mathbf{FF}_{1})\circ(\operatorname{id}+\mathbf{SA}_{1}),$ (5)
where $\mathbf{SA}_{\ell}$ and $\mathbf{FF}_{\ell}$ denote the causal attention and the feedforward layers at depth $\ell\in[L]$, respectively, and $\operatorname{id}$ denotes the identity function. Given an input sequence $x=(x_{1},\dots,x_{n})\in{\mathcal{V}}^{n}$, we define the initial sequence as: $f_{\mathrm{cot}}^{0}(x)\coloneq\mathbf{WE}(x)+\mathbf{PE}([n]),$ where the input tokens are first embedded via the word embedding function $\mathbf{WE}$ and the positional encoding $\mathbf{PE}$ Then, the continuous thought (Coconut) computes recursively as:
$f_{\mathrm{ct}}^{k+1}(x)\coloneq f_{\mathrm{ct}}^{k}(x)\cdot(\mathrm{TF}_{\mathrm{dec}}(f_{\mathrm{cot}}^{k}(x)+\mathbf{PE}(k))),$ (6)
where $\cdot$ denotes concatenation. The final output of the model after $T(n)$ steps is defined as the last output length $m$ tokens of $\mathrm{Dec}\,\left(\mathbf{OUT}\,\left(f_{\mathrm{ct}}^{T(n)}(x))\right)\right),$ where the final output is produced by a linear projection $\mathbf{OUT}$ and $\mathrm{Dec}(\cdot)$ is a decoding function that maps the output logits to a token in $\mathcal{V}$}: in the deterministic model, $\mathrm{Dec}(z)\coloneq\arg\max_{i\in[|{\mathcal{V}}|]}z_{i}$; in the stochastic model, $\mathrm{Dec}(z)\sim\mathrm{Multinomial}({z_{i}}/{\sum_{j}z_{j}})$, assuming $z_{i}>0$ for all $i$.
A.5 Looped Transformer
Definition A.4 (Looped TF).
Let the Transformer block be defined as the composition:
$\mathrm{TF}\coloneq(\operatorname{id}+\mathbf{FF}_{L})\circ(\operatorname{id}+\mathbf{SA}_{L})\circ\cdots\circ(\operatorname{id}+\mathbf{FF}_{1})\circ(\operatorname{id}+\mathbf{SA}_{1}),$ (7)
@@ A.5 Looped Transformer
$f_{\mathrm{loop}}^{k+1}(x)\coloneq\mathrm{TF}\left(f_{\mathrm{loop}}^{k}(x)\right).$ (8)
@@
Figure 8: The models of reasoning paradigms based on iterative use of Transformer models.
Appendix B Deferred Proofs for Section 3
B.1 Precision Modeling
We focus on signed floating-point numbers, following (Li et al., 2024), but omit exponents for simplicity.
Definition B.1 (Floating-point Representation, cf. (Li et al., 2024)).
@@ B.1 Precision Modeling
Definition B.2 (Correct Rounding, cf. (Li et al., 2024)).
@@
Definition B.3 (Summation with Iterative Rounding (Li et al., 2024)).
@@
Based on this definition, all computations in the Transformer block of Definition A.1 can be represented in finite precision. The inner product and the matrix product are defined as
@@
B.2 Definition of Assumption 3.4
Our definition of polynomially efficient approximation follows that of Feng et al. (2023), but differs in scope: while their framework targets real-valued functions, ours applies to symbolic functions.
Definition B.4 (Polynomially-efficient approximation).
We say that a function $f_{n}:\Sigma^{\ell(n)}\to\Sigma$ admits a polynomially efficient approximation if, for a sufficiently small error tolerance $0<\delta\leq\tfrac{1}{3}$, there exists a feedforward network $\mathrm{FF}:\mathbb{F}_{s(n)}^{\ell(n)\cdot|\Sigma|}\to\mathbb{F}_{s(n)}^{|\Sigma|},\ s(n)=O(\log n),$ such that the following holds: for every input ${\bm{x}}=(x_{1},\ldots,x_{\ell(n)})\in\Sigma^{\ell(n)}$,
@@ B.3 Technical lemmas
B.3.1 Orthogonal Vectors
We follow the notation of (Li et al., 2024). For any positive integer $s\in\mathbb{N}^{+}$ and $x\in\{0,1,\dots,2^{s}-1\}$, we denote by $\mathsf{bin}_{s}(x)\in\{0,1\}^{s}$ the standard binary representation of $x$ using $s$ bits, defined such that $x=\sum_{i=1}^{s}2^{i}\cdot(\mathsf{bin}_{s}(x))_{i}.$ We further define the signed binary encoding of $x$, denoted by $\mathsf{sbin}_{s}(x)\in\{-1,1\}^{s}$, as $\mathsf{sbin}_{s}(x)=2\cdot\mathsf{bin}_{s}(x)-(1,\ldots,1).$ Let ${\bm{x}},{\bm{y}}\in\mathbb{R}^{s}$ be two vectors of the same length. We define their interleaving, denoted by ${{\bm{x}}}^{\frown}{{\bm{y}}}\in\mathbb{R}^{2s}$, as follows: $({{\bm{x}}}^{\frown}{{\bm{y}}})_{2i-1}=x_{i},({{\bm{x}}}^{\frown}{{\bm{y}}})_{2i}={\bm{y}}_{i}$ for all $i\in[s].$ The orthogonal vectors under finite-precision arithmetic can be:
Lemma B.5 (Li et al., 2024).
@@ B.3.1 Orthogonal Vectors
B.3.2 Position Selector
Lemma B.6.
For any $m\in\mathbb{N}^{+}$ and ${\bm{x}}\in\mathbb{F}_{s}^{m}$ with $x_{i}>0$ for all $i\in[m]$, there exists a feedforward layer $\mathrm{FF}:\mathbb{F}_{s}^{2m}\to\mathbb{F}_{s}^{2m}$, for any $i\in[m]$, such that
@@ B.3.2 Position Selector
Let the input be ${\bm{z}}=({\bm{x}},{\bm{e}}_{i})\in\mathbb{F}_{s}^{2m}$. Set the weight ${\bm{W}}_{1}\in\mathbb{F}_{s}^{m\times 2m}$ and bias ${\bm{b}}\in\\ mathbb{F}_{s}^{m}$ by
${\bm{W}}_{1}=\bigl[\,{\bm{I}}_{m}\;\;\;-B_{s}{\bm{I}}_{m}\,\bigr],\qquad{\bm{b}}={\bm{0}},$ (14)
@@
Next, set the second linear layer ${\bm{W}}_{2}\in\mathbb{F}_{s}^{2m\times m}$ by
${\bm{W}}_{2}=\begin{bmatrix}-{\bm{I}}_{m}\\ \mathbf{0}_{m\times m}\end{bmatrix}$. (17)
@@
B.3.3 Feedforward layers
Lemma B.7.
@@ B.3.3 Feedforward layers
For each $i\in[N]$, by Definition B.4, there exist width $w_{i}\in\mathbb{N}$ and parameters
${\bm{W}}^{(i)}_{1}\in\mathbb{F}_{s}^{w_{i}\times\ell_{i}|\Sigma|},\quad{\bm{W}}^{(i)}_{2}\in\mathbb{F}_{s}^{|\Sigma|\times w_{i}},\quad{\bm{b}}^{(i)}\in\mathbb{F}_{s}^{w_{i}}$ (20)
@@
${\bm{W}}_{1}\coloneqq\bigoplus_{i=1}^{N}{\bm{W}}^{(i)}_{1},\qquad{\bm{W}}_{2}\coloneqq\bigoplus_{i=1}^{N}{\bm{W}}^{(i)}_{2},\qquad{\bm{b}}\coloneqq\bigoplus_{i=1}^{N}{\bm{b}}^{(i)}.$ (21)
@@
$\mathrm{FF}({\bm{x}})\coloneqq{\bm{W}}_{2}\mathrm{ReLU}({\bm{W}}_{1}{\bm{x}}+{\bm{b}})$ (22)
@@
Lemma B.8.
Let $f:\Sigma^{\ell}\to\Sigma$ be a function that admits polynomially-efficient approximation (Definition B.4). Then, there exist two feedforward layers
$\mathrm{FF}_{1}:\mathbb{F}_{s}^{(1+\ell)|\Sigma|+1}\to\mathbb{F}_{s}^{(1+\ell)|\Sigma|+1},\quad\mathrm{FF}_{2}:\mathbb{F}_{s}^{(1+\ell)|\Sigma|+1}\to\mathbb{F}_{s}^{(1+\ell)|\Sigma|+1},$ (23)
@@
Let $n\coloneqq|\Sigma|$ and $L\coloneqq\ell n$. Write ${\bm{u}}=(\bm{e}(x_{1}),\ldots,\bm{e}(x_{\ell}))\in\{0,1\}^{L}$. By Definition B.4, there exist $w_{f}\in\mathbb{N}$, matrices ${\bm{W}}^{(f)}_{1}\in\mathbb{F}_{s}^{w_{f}\times L}$, ${\bm{W}}^{(f)}_{2}\in\mathbb{F}_{s}^{n\times w_{f}}$, and bias ${\bm{b}}^{(f)}\in\mathbb{F}_{s}^{w_{f}}$ such that
@@
${\bm{W}}^{(1)}_{1}=\begin{bmatrix}{\bm{0}}&{\bm{W}}^{(f)}_{1}&\mathbf{0}_{\,w_{f}\times 1}\\ {\bm{0}}&{\bm{I}}_{L}&\mathbf{0}_{\,L\times 1}\end{bmatrix},\quad{\bm{b}}^{(1)}=\begin{bmatrix}{\bm{b}}^{(f)}\\ {\bm{0}}_{L}\end{bmatrix},\quad{\bm{W}}^{(1)}_{2}=\begin{bmatrix}{\bm{W}}^{(f)}_{2}&\,{\bm{0}}\\ \mathbf{0}_{\,L\times w_{f}}&\,{\bm{0}}\\ \mathbf{0}_{\,1\times w_{f}}&{\bm{0}}\end{bmatrix}$, (26)
and define $\mathrm{FF}_{1}({\bm{0}},{\bm{u}},t)\coloneqq{\bm{W}}^{(1)}_{2}\,\mathrm{ReLU}\big({\bm{W}}^{(1)}_{1}({\bm{0}},{\bm{u}},t)+{\bm{b}}^{(1)}\big)$. Then, it holds that
@@
Thus we have $(\operatorname{id}+\mathrm{FF}_{1})({\bm{0}},{\bm{u}},t)=\bigl(\mathrm{FF}_{f}({\bm{u}}),\,{\bm{u}},\,t\bigr).$
@@
${\bm{W}}^{(2)}_{1}=\begin{bmatrix}2{\bm{I}}_{n}&\mathbf{0}_{n\times L}&M{\bm{1}}_{n}\\ 2{\bm{I}}_{n}&\mathbf{0}_{n\times L}&M{\bm{1}}_{n}\\ {\bm{I}}_{n}&\mathbf{0}_{n\times L}&\mathbf{0}_{n\times 1}\\ -{\bm{I}}_{n}&\mathbf{0}_{n\times L}&\mathbf{0}_{n\times 1}\end{bmatrix},\ {\bm{b}}^{(2)}=\begin{bmatrix}-M{\bm{1}}_{n}\\ (1-M){\bm{1}}_{n}\\ {\bm{0}}_{n}\\ {\bm{0}}_{n}\end{bmatrix},\ {\bm{W}}^{(2)}_{2}=\begin{bmatrix}\ {\bm{I}}_{n}&-{\bm{I}}_{n}&-{\bm{I}}_{n}&\ {\bm{I}}_{n}\ \\ \ \mathbf{0}_{\,L\times n}&\mathbf{0}_{\,L\times n}&\mathbf{0}_{\,L\times n}&\mathbf{0}_{\,L\times n}\\ \ \mathbf{0}_{\,1\times n}&\mathbf{0}_{\,1\times n}&\mathbf{0}_{\,1\times n}&\mathbf{0}_{\,1\times n}\end{bmatrix}$. (30)
@@
$\mathrm{FF}_{2}({\bm{z}},{\bm{u}},t)$ $={\bm{W}}^{(2)}_{2}\,\begin{bmatrix}\mathrm{ReLU}\bigl(2{\bm{z}}+Mt{\bm{1}}_{n}-M{\bm{1}}_{n}\bigr)\\ \mathrm{ReLU}\bigl(2{\bm{z}}+Mt{\bm{1}}_{n}+(1-M){\bm{1}}_{n}\bigr)\\ \mathrm{ReLU}({\bm{z}})\\ \mathrm{ReLU}(-{\bm{z}})\end{bmatrix}$ (31)
@@
Therefore, the composition satisfies $(\operatorname{id}+\mathrm{FF}_{2})\circ(\operatorname{id}+\mathrm{FF}_{1})\bigl({\bm{0}},{\bm{u}},t\bigr)=(t\cdot\bm{e}(f({\bm{x}})),\ {\bm{u}},\ t).$
B.4 Proof for Theorem 3.5
@@
$\bm{\mathrm{sbinpred}}_{s(n)}(v_{i})\coloneq\left(\mathsf{sbin}_{s(n)}(\mathrm{pred}(v_{i})_{0}),\,\ldots,\,\mathsf{sbin}_{s(n)}(\mathrm{pred}(v_{i})_{C_{\mathrm{max}}(n)})\right).$ (37)
@@
${\bm{q}}_{i,h}$ $={\mathsf{sbin}_{s(n)}(\mathrm{pred}(v_{i+1})_{h})}^{\frown}{1_{s(n)}},$ (39)
@@
where vectors of different lengths are zero-padded to match the dimension. By Lemma B.5, each attention head of the last position $i=n+k$ retrieves the predecessor’s value of $v_{n+k}$
${\bm{a}}_{n+k,h}=\bm{e}\bigl(v_{\mathrm{pred}(v_{n+k+1})_{h}}(x)\bigr)$ (42)
@@
${\bm{O}}({\bm{a}}_{i,1},\ldots,{\bm{a}}_{i,H})=({\bm{0}},\,\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{0}}(x)),\,\ldots,\,\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{C_{\mathrm{max}(n)}}}(x)),\,{\bm{0}}),$ (43)
@@
$\underbrace{\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{0}}(x)),\,\ldots,\,\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{C_{\mathrm{max}(n)}}}(x))}_{\text{updated}}\,,{\bm{0}}_{C_{\mathrm{sum}}(n)|\Sigma|}\Big).$ (45)
@@
By Lemma B.7, a single feed-forward layer can approximate multiple functions by partitioning the input into blocks. The first feed-forward layer then places the arguments, gathered by attention, into the correct positions. By Lemma B.6, where the vector ${\bm{e}}_{i}$ therein corresponds to ${\bm{1}}_{|\Sigma|}$ here, the hidden state at the last position, denoted by ${\bm{h}}^{(1)}_{n+k}$, becomes
$\Big(({\bm{h}}^{(0.5)}_{n+k})_{1:r},\big\|_{j=1}^{|\mathcal{F}|}\,\big(\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{1}}(x))\cdot 1,\ldots,\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{C_{j}(n)}}(x))\cdot 1\Big),$ (46)
@@
By Assumption 3.4 and Lemmas B.8 and B.7, there exist feed-forward layers $\mathrm{FF}_{2},\mathrm{FF}_{3}:\mathbb{F}_{s(n)}^{m}\to\mathbb{F}_{s(n)}^{m}$ such that, for every input tuple ${\bm{x}}_{j}\coloneq(x^{(j)}_{1},\ldots,x^{(j)}_{C_{j}(n)})\in\Sigma^{C_{j}(n)}$ for $j\in[|\mathcal{F}|]$ and every ${\bm{t}}\in\{0,1\}^{|\mathcal{F}|}$, the composition $\mathcal{FF}\coloneq(\operatorname{id}+\mathrm{FF}_{3})\circ(\operatorname{id}+\mathrm{FF}_{2})$ satisfies
@@
$=\mathcal{FF}\left(*,\,\bm{e}(v_{n+k+1}),\,*,\,\bigg\|_{j=1}^{|\mathcal{F}|}\,\big(\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{1}}(x)),\ldots,\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{C_{j}(n)}}(x))\big)\right)$ (49)
$=\Biggl(\sum_{j=1}^{|\mathcal{F}|}\bm{e}(v_{n+k+1})_{j}\cdot\bm{e}\Big(f_{j}\big(\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{1}}(x)),\ldots,\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{C_{j}(n)}}(x))\big)\Big),\;*\Biggr)$ (50)
$=\Biggl(\bm{e}\Big(f_{l}\big(\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{1}}(x)),\ldots,\bm{e}(v_{\mathrm{pred}(v_{n+k+1})_{C_{l}(n)}}(x))\big)\Big),\;*\Biggr),\,\text{where}\quad\bm{e}(v_{n+k+1})={\bm{e}}_{l}$ (51)
@@
B.5 Proof of Theorem 3.6 for Looped Transformer
Figure 9: Illustration of the role of the attention and feedforward layers in looped TFs for evaluating DAGs: the attention layer uniformly attends to and aggregates all inputs at each position, while the feedforward layer simultaneously simulates the functions of the nodes.
@@
In the proof, we assume that the computation graph contains at most $n$ output nodes. This assumption is without loss of generality: if the number of output nodes exceeds the number of input nodes, we can simply pad the input with dummy nodes (e.g., fixed zeros), thereby reducing the setting to the same case.
We construct a model in which (1) the attention layer aggregates the inputs, and (2) the looped feed-forward layer performs the computation of all nodes in parallel, as illustrated in Figure 9. We first show that a feedforward layer followed by an attention layer can copy all input tokens to each position. Assume, given an input sequence $x=(x_{1},x_{2},\ldots,x_{n})\in\Sigma^{n}$ and one-hot encoding $\bm{e}:\Sigma\to\{0,1\}^{|\Sigma|}$. Assume each input at position $i\in[n]$ is embedded as